# From Activation to Initialization: Scaling Insights for Optimizing Neural Fields

### v1

@@ 3 Notation
where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise. The notation $[m]$ is defined by $[m]=\{1,\ldots,m\}$. For a weight matrix $W_{k}$ at layer k, the notation $W_{k}^{0}$ will denote the initialization of that weight matrix. These are the initial weights of the network before training begins under a gradient descent algorithm. The networks we use in the paper will all be trained using the MSE loss function, which we denote by $\mathcal{L}$. Our theoretical results will be primarily for the case where $\phi$ is given by one of the activation sine, sinc, Gaussian or wavelet. We remind the reader that the sinc function is defined by: $sinc(x)=\frac{sin(x)}{x}$ for $x\neq 0$ and $sinc(0)=1$. For more details on how these activations are used within the context of neural fields we refer the reader to [29, 23, 26, 24].
We will always assume the data set $X$ consists of i.i.d sub-gaussian vectors and for the theoretical proofs will assume they are normalized to have norm $||X_{i}||_{2}=1$. More details on data assumptions can be found in sec. 1 of the supp. material.
@@ 4.3 Analyzing the proof methodology
$\sigma_{0}^{2}\geq 16\sqrt{N}\sqrt{n_{1}}\sqrt{2\mathcal{L}(\theta_{0})}||W_{2}^{0}||_{2}$ (4.6)
@@
$||W||_{2}=\mathcal{O}(\sqrt{n_{2}}/\sqrt{n_{1}}).$ (4.9)
@@ 4.4 Designing new initializations
As suggested by the discussion in sec. 4.3 the variance of the Gaussian we sample from for the last layer is smaller by a factor of $1/\sqrt{n_{L-1}}$. The biases of the network will be initialized to 00 or a very small number such as $0.01$ as is the standard practice for many common normal initializations. Fig. 2, gives a diagrammatic rendition of initialization 1.
@@
where $\mathcal{U}(a,b)$ denotes the Uniform distribution on the interval $[a,b]$. The biases can be initialized to be 00 or a very small number. In sec. 1 of the supp. material we give another way to initialize the biases.
@@ 5.1 Practical Validation of the Theoretical Analysis
In our investigation, we performed a 1-dimensional curve fitting experiment on the function $f(x)=\sin(2\pi x)+\sin(6\pi x)+\sin(10\pi x)$. We systematically sampled the curve at intervals of $10$, $20$, $50$, $75$, $100$, $125$, $150$, $175$, and $200$ points, creating nine datasets of varying sizes for our training data.
@@ 5.2 Single Image Super Resolution
We test our initializations on an image super resolution task. We take the approach considered in [26], where a 4×4$\times$4 × super resolution is conducted on an image from the DIV2K dataset [1, 34]. The problem is cast as solving $y=Ax$, where the operator $A$ implements a 4×4$\times$4 × downsampling (with no aliasing). We then solve for $x$ as the output of a neural field.
@@
Figure 5: The figure shows the results for a 4×4$\times$4 × single image super resolution with four normal initializations and four uniform initializations. Networks initialized with initialization 1 (our) and initialization 2 (our) produced the highest train dB and SSIM at convergence. Zoom in for better viewing.
@@ 5.3 Occupancy Fields
We optimize a binary occupancy field, which represents a 3D shape as the decision boundary of a neural network  [35, 12]. We use the thai statue instance obtained from XYZ RGB Inc. We trained two groups of four networks, each activated with the Gabor wavelet [26]. The first group of four networks were trained with four normal initializations, Lecun normal, Xavier normal, Kaiming normal and initialization 1 (see (4.10), (4.11)). The second group of four networks were trained with LeCun uniform, Xavier uniform, Kaiming Uniform and initialization 2 (see (4.15), (4.16)). All networks were trained with the Adam optimizer. For accuracy testing we used the performance metric given by Intersection Over Union (IOU) . Fig. 6 and table 1 details the results of the experiments. As can be seen from the figure, both our initializations converge to a higher PSNR, $2-4$dB higher than the others, and have the highest IOU. Figures of reconstructed meshes can be found in the sec. 2 of supp. material.
@@ 5.5 Physics Informed Neural Networks (PINNs)
where $u(x,y,t)$ denotes the $x$-component of the velocity field of the fluid, and $v(x,y,t)$ denotes the $y$-component of the velocity field. The term $p(t,x,y)$ is the pressure. The domain of the problem is $[-15,25]\times[-8,8]\times[0,20]$.

### v2

@@ 3 Notation
where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise. The notation $[m]$ is defined by $m]=\{1,\ldots,m\}$. For a weight matrix $W_{k}$ at layer k, the notation $W_{k}^{0}$ will denote the initialization of that weight matrix. These are the initial weights of the network before training begins under a gradient descent algorithm. The networks we use in the paper will all be trained using the MSE loss function, which we denote by $\mathcal{L}$. Our theoretical results will be primarily for the case where $\phi$ is given by one of the activation sine, sinc, Gaussian or wavelet. We remind the reader that the sinc function is defined by: $sinc(x)=\frac{sin(x)}{x}$ for $x\neq 0$ and $sinc(0)=1$. For more details on how these activations are used within the context of neural fields we refer the reader to [28, 23, 26, 24].
We will always assume the data set $X$ consists of i.i.d sub-gaussian vectors and for the theoretical proofs will assume they are normalized to have norm $X_{i}||_{2}=1$. More details on data assumptions can be found in sec. 1 of the supp. material.
@@ 4.3 Analyzing the proof methodology
[…]
@@
$W||_{2}=\mathcal{O}(\sqrt{n_{2}}/\sqrt{n_{1}}).$ (4.9)
@@ 4.4 Designing new initializations
As suggested by the discussion in sec. 4.3 the variance of the Gaussian we sample from for the last layer is smaller by a factor of $1/\sqrt{n_{L-1}}$. The biases of the network will be initialized to 0 or a very small number such as $0.01$ as is the standard practice for many common normal initializations. Fig. 2, gives a diagrammatic rendition of initialization 1.
@@
where $\mathcal{U}(a,b)$ denotes the Uniform distribution on the interval $[a,b]$. The biases can be initialized to be 0 or a very small number. In sec. 1 of the supp. material we give another way to initialize the biases.
@@ 5.1 Practical Validation of the Theoretical Analysis
In our investigation, we performed a 1-dimensional curve fitting experiment on the function $f(x)=\sin(2\pi x)+\sin(6\pi x)+\sin(10\pi x)$. We systematically sampled the curve at intervals of $10$, 2020, $50$, $75$, $100$, $125$, $150$, $175$, and $200$ points, creating nine datasets of varying sizes for our training data.
@@ 5.2 Single Image Super Resolution
We test our initializations on an image super resolution task. We take the approach considered in [26], where a $4\times$ super resolution is conducted on an image from the DIV2K dataset [1, 34]. The problem is cast as solving $y=Ax$, where the operator $A$ implements a $4\times$ downsampling (with no aliasing). We then solve for $x$ as the output of a neural field.
@@
Figure 5: The figure shows the results for a $4\times$ single image super resolution with four normal initializations and four uniform initializations. Networks initialized with initialization 1 (our) and initialization 2 (our) produced the highest train dB and SSIM at convergence. Zoom in for better viewing.
@@ 5.3 Occupancy Fields
We optimize a binary occupancy field, which represents a 3D shape as the decision boundary of a neural network  [35, 12]. We use the thai statue instance obtained from XYZ RGB Inc. We trained two groups of four networks, each activated with the Gabor wavelet [26]. The first group of four networks were trained with four normal initializations, Lecun normal, Xavier normal, Kaiming normal and initialization 1 (see (4.10), (4.11)). The second group of four networks were trained with LeCun uniform, Xavier uniform, Kaiming Uniform and initialization 2 (see (4.15), (4.16)). All networks were trained with the Adam optimizer. For accuracy testing we used the performance metric given by Intersection Over Union (IOU) . Fig. 6 and table 1 details the results of the experiments. As can be seen from the figure, both our initializations converge to a higher PSNR, 242-4dB higher than the others, and have the highest IOU. Figures of reconstructed meshes can be found in the sec. 2 of supp. material.
@@ 5.5 Physics Informed Neural Networks (PINNs)
where $u(x,y,t)$ denotes the $x$-component of the velocity field of the fluid, and $v(x,y,t)$ denotes the $y$-component of the velocity field. The term $p(t,x,y)$ is the pressure. The domain of the problem is $15,25]\times[-8,8]\times[0,20]$.
@@ 7 Limitations
8 Acknowledgements
Simon Lucey acknowledges support from the Australian Research Council (ARC) through the Discovery Project DP220103803.